OpenStudy (loser66):

Suppose A\(\in L(V)\) such that rank (A)=1. Prove that there exists \(\vec u_0 \in V\) and y\(\in V'\) such that \(A\vec v=[\vec v, y]\vec u_0\) for every \(\vec v \in V\). Find all eigenvalues of A and indicate their multiplicities. What is Tr(A)? Appreciate any help. Please

OpenStudy (loser66):

@zzr0ck3r

OpenStudy (zzr0ck3r):

its been a while since ive done real linear algebra

OpenStudy (zzr0ck3r):

what is L?

OpenStudy (loser66):

it's just linear transformation set

OpenStudy (zzr0ck3r):

Is V' V transpose?

OpenStudy (loser66):

V is field, and V' is dual of V. V' is set of operator on V

OpenStudy (zzr0ck3r):

yeah I cant help you:( I have only taken one linear algebra class, so I really only know basic stuff. This is why I can never help your posts

OpenStudy (loser66):

@wio

Miracrown (miracrown):

Since rank(A)=1, we see that the dimension of the linear transformation l_A with matrix A is equal to dim(V)-1. That means the dimension of the image of l_A is equal to 1. Let L_A: V->V be the linear transformation with matrix A. Since rank(A)=1, we have that dim(Im(l_A))=1, where Im(l_A) is the image of l_A. Then Im(l_A) is spanned by a vector u_0 in V. That is, Im(l_A)=<u_0>. For any v in V, l_A(v) is in Im(l_A). So we have that l_A(v)=k*u_0 for some scalar k. It is clear that l_A(v)=Av by definition. Therefore, l_A(v)=Av=[v,y]u_0 for all v in V, where [v,y]=k According to Riesz representation theorem, we have that l_A(v)=Av=[v,y]u_0 for some y.

Miracrown (miracrown):

That is the proof for the first part. All of the eigenvalues of A are 0 and 1. the multiplicity of 0 is dim(V)-1, and the multiplicity of 1 is 1.

Miracrown (miracrown):

The eigenvalues of A are 0 and [u_0, y]. The multiplicity of 0 is dim(V)-1. The multiplicity of [u_0, y] is 1. Trace(A)=n*0+1*[u_0,y]=[u_0,y]. (n-1)*0 instead of n*0. Tr(A)=Trace(A).